OpenStudy (anonymous):

Find values of a and b such that the function f(x)=ax+b /(x^2)−1 has a local extreme value of 1 at x=3. Is this extreme value a local maximum, or a local minimum? Give reasons for your answer.

OpenStudy (freckles):

so first let's try to find f'

OpenStudy (anonymous):

will do.

OpenStudy (freckles):

also is it (ax+b)/(x^2-1)?

OpenStudy (anonymous):

yup

OpenStudy (zale101):

Find the derivative and set it equal to zero. After solving for x, those will be the critical points. Do you know how to take the derivative of this function?

OpenStudy (anonymous):

\[\frac{ ax+b }{ x^2 -1 }\] = f(x)

OpenStudy (anonymous):

that's the original, and to get the derivative i use quotient rule.

OpenStudy (anonymous):

So i think it's : \[\frac{ -ax^2 -a -2bx }{ (x^2 -1)^2 }\]

OpenStudy (freckles):

you will have a system of equations to solve by the way \[f(3)=1 \\ \text{ then set the numerator of } f' =0 \text{ then solve for x}\\ \text{ then use that that x is suppose to be 3}\]

OpenStudy (freckles):

himm I think I got a little bit of a different numerator

OpenStudy (freckles):

oh you combined like terms lol

OpenStudy (freckles):

sounds good

OpenStudy (freckles):

\[-ax^2-2bx-a=0\] solve for x

OpenStudy (freckles):

use the quadratic formula

OpenStudy (anonymous):

okay let's see...

OpenStudy (anonymous):

I got: \[x = a + or - {\sqrt{a^2-8abx}}\]

OpenStudy (anonymous):

it's supposed to be above -2a

OpenStudy (anonymous):

\[\frac{ a + or - \sqrt {a^2 - 8abx} }{ -2a}\]

OpenStudy (freckles):

\[-ax^2-2bx-a=0 \\ x=\frac{-(-2b) \pm \sqrt{(-2b)^2-4(-a)(-a)}}{2(-a)} \\ x=\frac{2b \pm \sqrt{4b^2-4a^2}}{-2a}\\ x=\frac{b \pm \sqrt{b^2-a^2}}{-a} \\ \text{ we want one of these x's to be 3 }\] so we have \[\frac{b \pm \sqrt{b^2-a^2}}{-a}=3 \] and we also have the other equation: from f(3)=1 \[f(3)=\frac{3a+b}{9-1}=\frac{3a+b}{8}=1 \\ 3a+b=8 \\ b=8-3a\] so i would plug this b in the ther other equation

OpenStudy (freckles):

and hopefully we can solve for a

OpenStudy (anonymous):

oh i think i was using quadtratic on the wrong thing no clue where the x came from

OpenStudy (freckles):

and don't worry the a's do not look that bad i got integers

OpenStudy (anonymous):

so i plug b into the quadtratic formula answer?

OpenStudy (freckles):

yeah replace the b's with (8-3a)

OpenStudy (freckles):

having trouble?

OpenStudy (anonymous):

it was a small mistake during my calculations again. sorry

OpenStudy (anonymous):

 so like this?